Evaluate the expression (9-9i)(6+4i) and write the result in the form a+ b i.
Answer by nyc_function(2733)

(9-9i)(6+4i) =


54 + 36i - 54i -36(i^2) =


54 + 36i - 54i -36(-1) =


54 + 36i - 54i + 36 =


90 - 18i